OpenStudy (anonymous):

There are only two forces on an object that has a mass of 10.0 kg. Both forces have magnitude of 10.0 N. The angle between the forces is 60 degrees. What is the magnitude of the object's acceleration?

OpenStudy (***[isuru]***):

Hi, When there are 2 forces of equal magnitude acting with an angle of 60 degree to each other the Resultant force is 10√3 N ( this is calculated using R^2 = P^2 + Q^2 + 2PQcos(x) ) using F = ma 10√3 = 10 * a a = √3 m/s^2